Solve this equation by the substitution method. Help
{2y=x+2
{6x-12y=0

Found 2 solutions by robertb, stanbon:
Answer by robertb(5830)

From the top equation, x = 2y-2. Substitute this into the bottom equation:6(2y-2)-12y = 0.
12y-12-12y = 0.
-12 = 0. This obviously a false statement. Thus the system has no solution set, it is inconsistent.

Answer by stanbon(75887)

Solve this equation by the substitution method. Help
{2y=x+2
{6x-12y=0
---------------
solve the 1st eq. for "x":
x = 2y-2
---
Substitute into the 2nd eq. and solve for "y":
6(2y-2)-12y = 0
12y-12-12y = 0
-12 = 0
=============
That is a contradiction.
There is no solution for the system of equations.
The lines are parallel; they do not meet at some (x,y) point.
The system is said to be inconsistent.
======================
Cheers,
Stan H.
============